# A Survey on the Application of Large Language Models in Scenario-Based Testing of Automated Driving Systems

### v1

@@ III-B Scenario Abstraction Level
According to [37, 38], scenarios are organized into four levels of abstraction: 1) Functional scenarios represent high-level traffic situations described in natural language using domain-specific terms. 2) Abstract scenarios are defined by transforming natural language descriptions into formats suitable for machine interpretation, often supported by ontologies. 3) Logical scenarios define the parameter ranges of the state values used for scenario representation. 4) Concrete scenarios use specific state values to ensure their reproducibility and to enable test methods to execute the scenario.
IV LLM applications by phase
for tree=
grow=east, parent anchor=east,
child anchor=west,
edge=ieeecyan, line width=1pt,
edge path=
[…]
rounded corners, draw, align=center, font=,
minimum size=3mm,
for children=
[Section IV LLM Applications by Phase, fill=ieeecyan!50, draw=none[IV-E ADS Assessment, fill=ieeecyan!50, name=ADSAssessment, draw=none
Safety performance evaluation [39, 40]
Intelligence level evaluation [41]
[IV-D Test Execution, fill=ieeecyan!50, name=TestExecution, draw=none
[Scenario optimization, fill=ieeecyan!50, draw=none
Syntax correction [42, 43, 44, 45]
Fidelity refinement [46]
[Simulation setup automation, fill=ieeecyan!50, draw=none
[39, 47, 48, 49]
[Anomaly detection, fill=ieeecyan!50, draw=none
[50]
[IV-C Scenario Selection, fill=ieeecyan!50, name=ScenarioSelection, draw=none
[Realism assessment, fill=ieeecyan!50, draw=none
[51, 52]
[IV-B Scenario Generation, fill=ieeecyan!50, name=ScenarioGeneration, draw=none
[LLM as an executable scenario generator, fill=ieeecyan!50, draw=none
Template filling [44]
End-to-end generation [43, 46, 53, 54]
Hybrid generation [42, 47, 48, 49, 55, 56, 57, 58]
[LLM as a standardized format generator, fill=ieeecyan!50, draw=none
OpenSCENARIO [55]
AVUnit [59]
[LLM as an intermediate format generator, fill=ieeecyan!50, draw=none
Driving policies [60, 61, 62]
Scenario elements [63, 64, 65, 66]
Functional scenario [67]
Abstract scenario [68]
Logical scenario [69]
[LLM as a data interpreter, fill=ieeecyan!50, draw=none
Accident report [70]
Domain-specific knowledge [71]
Naturalistic driving data [72, 73, 74, 75]
[LLM as a human-machine interface, fill=ieeecyan!50, draw=none
Structured scenario representation [76, 77, 39, 78, 79, 80]
Loss function [81]
Executable code [82, 83, 84]
[IV-A Scenario Source, fill=ieeecyan!50, name=ScenarioSource, draw=none
[Data retrieval, fill=ieeecyan!50, draw=none
Video [85]
Image [86, 87, 88]
[Data labeling, fill=ieeecyan!50, draw=none
[89]
[Data enrichment
through harzard analysis, fill=ieeecyan!50, draw=none
HARA [90, 91, 92]
STPA [93, 94, 95]
[Data enrichment, fill=ieeecyan!50, draw=none
Driving trajectory [96, 97, 98]
Future driving video [99]
Photo-realistic driving scene [100]
The scenario-based testing workflow, as illustrated in Figure 2, is initiated with scenario sources, which may include synthesized data, naturalistic driving measurements, or human expert knowledge. These sources are utilized to generate test scenarios, which are subsequently organized and stored within a scenario database. Scenario selection is then performed according to predefined criteria, followed by execution on simulation platforms such as CARLA [101] to evaluate the performance of ADSs. Finally, test results are analyzed and reported in the ADS assessment phase.
In this section, the application of LLMs in scenario-based testing within different testing phases are examined, along with corresponding implementation details. An overview of the research structure for this section is presented in Figure 3.
IV-A Scenario Source
Scenario-based testing of ADSs heavily relies on data, which serve as the basis for generating test scenarios. As a result, considerable attention has been directed toward scenario sources to support this process. As noted in [15, 19], these source can be classified into data-based or knowledge-based categories. Data may be obtained from real-world driving environments or synthetically generated using real-world contexts and expert input. Surveyed literature indicates that LLMs are primarily employed for tasks such as data enrichment, labeling, and retrieval. The implementation strategies associated with these applications are described below.
IV-A1 Data Enrichment
The acquisition of real-world driving data often falls short of meeting the evolving demands of ADSs, primarily due to the high costs and time-intensive nature of data collection and annotation. As a result, data synthesis has become an essential component for ADS testing [102]. LLMs have been employed to generate driving trajectories based on natural language instructions. In [96], LLMs are utilized for both trajectory-conditioned language generation and language-conditioned trajectory synthesis. In [97], right-hand driving trajectories are synthesized by LLMs using left-hand driving data as a reference. In [98], LLM is used to translate brief descriptions of vehicle interactions into realistic driving behaviors and trajectories by incorporating human-like driving logic.
Jia et al. [99] further expand data enrichment by combining an LLM with a video diffusion model to generate future driving scenes in video format. This enables trajectory generation and visual simulation based on natural language–driven vision–action pairs. Furthermore, LLMs have been adopted for ground truth generation; for example, Wei et al. [100] propose ChatSim, a framework that enables the creation of editable, photo-realistic 3D driving scenes through natural language commands. Within this framework, LLMs are used to facilitate user interaction and to enable the integration of external digital assets.
IV-A2 Data Enrichment Through Hazard Analysis
Hazard Analysis and Risk Assessment (HARA) and Systems Theoretic Process Analysis (STPA) are originally developed for functional safety and system-level hazard analysis, respectively. These methods are also commonly applied within the scenario-based approach described in ISO 21448 (cf. [103, 104, 16]) to systematically identify hazardous scenarios, gain insights into potential hazards (cf. [105, 106, 107]) and guide the generation of high-risk scenario seeds (cf. [108, 109, 110]).
HARA automation using LLMs has been explored to improve the efficiency and accuracy of safety analysis in ADSs. At Qualcomm [90], LLMs enable the identification of 20% more hazardous scenarios in an Autonomous Emergency Braking (AEB) case study compared to manual methods. Volvo Cars researchers investigate LLMs to generate scenarios and malfunctions from functional descriptions using a general framework, without relying on standardized databases [91], thereby demonstrating their potential for domain-independent safety analysis. A follow-up study [92] applies LLMs to generate safety requirements and detect inconsistencies, which enhances efficiency in safety engineering while highlighting the need for expert validation to ensure accuracy.
STPA automation using LLMs has also been investigated. Charalampidou et al. [93] used ChatGPT-4 to generate loss scenarios for an unmanned aerial vehicle rescue system, although the results require human validation. Similarly, Diemert et al. [94] introduce a cooperative framework where LLMs assist in identifying hazards, though expert oversight is required due to inaccuracies. Qi et al. [95] extended this to automotive and energy systems, assessing human-LLM collaboration schemes.
IV-A3 Data Labeling
Manually annotate data is both costly and labor-intensive, thereby necessitating automated solutions to enhance the scalability and efficiency of dataset development [102]. To address this challenge, LLMs have been utilized to partially substitute human effort in the annotation process. Chen et al. [89] introduce the first benchmark for evaluating Large Vision-Language Models (LVLM) in the context of self-driving corner cases. In their work, LLMs are employed to automatically generate annotations in JSON111https://www.json.org/json-en.html format, which were subsequently verified through manual inspection to ensure data accuracy and quality.
IV-A4 Data Retrieval
Effective scenario generation requires the retrieval of relevant information from large-scale, heterogeneous datasets, which often include multimodal data in varying formats and structures. This inherent complexity presents challenges for unified processing and has traditionally been managed through rule-based methods, such as Structured Query Language (SQL) queries (cf. [111]), which demands significant manual effort. Recent studies have instead employed LLMs to enable more intuitive and flexible data retrieval, thereby reducing the reliance on predefined rules and manual querying.
Video retrieval efficiency has been demonstrated to improve through the application of LLMs. In the study by Knapp et al. [85], MLLMs are integrated with a vector database to enable natural language querying of driving logs. Scenario descriptions were generated from sensor data and video content, allowing users to navigate large-scale datasets without relying on SQL-based queries.
Image retrieval efficiency has also been improved through the application of LLMs. In [86], three LVLMs are employed to perform image retrieval based on natural language queries. Their performance is evaluated using precision-based metrics on a benchmark dataset containing diverse driving scenarios with varying scene complexity, weather conditions, and traffic levels. Additionally, Rigoll et al. [87] propose an annotation-free, object-level image retrieval approach for AD datasets by combining panoptic segmentation with a VLM to support natural language queries. While this approach enhances data accessibility without manual labeling, it is limited to single-object recognition and performs poorly in complex, dynamic scenes involving multiple objects. To overcome these limitations, Tang et al. [88] incorporate LLMs and Bird’s Eye View (BEV) representations to enhance text-to-scene retrieval. This approach is evaluated on nuScenes-Retrieval dataset, an extension of nuScenes dataset [112] with enriched and diverse textual annotations. Experimental results show that it achieves state-of-the-art performance, clearly outperforming existing baseline methods.
IV-B Scenario Generation
This section reviews previous studies that have employed LLMs for scenario generation. To enable execution within simulation platforms such as CARLA [101], the generated scenarios must be converted into standardized formats, such as ASAM OpenSCENARIO (XOSC) [113]. Given the inherent complexity of scenario generation, the process is typically decomposed into several sub-steps. When LLMs are utilized, their contributions are generally restricted to specific stages of the workflow rather than spanning the entire process. An overview of the functional roles played by LLMs throughout the scenario generation pipelines is illustrated in Figure 4, and the subsequent discussion is organized accordingly.
IV-B1 LLM as a Human-Machine Interface
In this subsection, LLMs are employed to translate natural language inputs provided by the user into structured information that facilitates downstream scenario generation, as illustrated in step ① of Figure 4. Specifically, LLMs have been used to interpret user input into structured scenario representations [76, 77, 39, 78, 79, 80], loss function formulations [81], and executable code [82, 83, 84]. The primary strategies adopted in these studies are schematically summarized in Figure 5.
Structured scenario representations derived from user inputs have been employed in various studies to support scenario generation. In [76], such structured information is employed to create executable files through a diffusion-based model. Similarly, in [77] and [39], it is used to trigger Python scripts for automated file generation. Building upon [77], Zhou et al. [114] evaluate the performance of six different LLMs in interpreting motorway functional scenarios, thereby offering a benchmark for model selection. In [78], structured representations are used as inputs to a transformer model for generating vehicle trajectories. In [79], LLMs convert natural language inputs into scenario elements, which are subsequently transformed into simulation-executable files. A unified diffusion-based framework is introduced in [80], where LLMs generate Proto222https://protobuf.dev/ constraints from user input to enable language-driven control of scene initialization and closed-loop simulation.
Loss functions and executable code have also been generated from natural language inputs using LLMs. In [81], a loss function is derived from user-provided descriptions and integrated into a diffusion model to facilitate executable file generation. Code generation from natural language inputs is further demonstrated in [82, 83]. In [82], the generated code, which includes interaction, vehicle, and map modules, is utilized by a transformer to synthesize vehicle trajectories. In [83], the generated code is applied to call functional libraries for producing vehicle trajectory data. In a related study, Nguyen et al. [84] employ a LLM to parse user language into driving behavior code, which is then used to guide reinforcement learning for generating diverse synthetic driving scenarios.
IV-B2 LLM as a Data Interpreter
This subsection reviews studies in which LLMs have been employed to interpret various data sources into natural language representations. This process is illustrated in step ② of Figure 4. Scenario-relevant information has been extracted from diverse inputs, including accident reports [70], domain-specific knowledge [71], and naturalistic driving data [72, 73, 74].
Accident reports have been interpreted by LLMs to support scenario reconstruction. In the study by Guo et al. [70], a framework is proposed in which textual accident reports are parsed into a structured JSON format, which is subsequently processed by a constraint solver to generate waypoints.
Additionally, domain-specific knowledge has been extracted using LLMs to aid scenario generation. Tang et al. [71] introduce a method in which technical documents are interpreted and transformed into structured ontological elements, thereby enabling automated ontology construction.
Naturalistic driving data have also been utilized to support sceario generation through LLM-based interpretation. Mei et al. [72] employ LLMs to identify the most threatening attacker in each scene from the Waymo Open Dataset [115], enabling the creation of high-risk scenarios for improving the testing and training of ADSs. In [73], LLMs are applied to interpret BEV maps derived from the nuPlan dataset [116], with the extracted spatial information subsequently used to guide a diffusion model for trajectory generation. Tian et al. [74] integrate reinforcement learning and LLMs to optimize safety-critical driving scenarios within the HighwayEnv [117] simulation environment. Extending these efforts into real-time application, Mei et al. [75] focus on online scenario generation by using retrieval-augmented LLMs to infer dangerous driving behaviors and synthesize adversarial trajectories based on historical states.
IV-B3 LLM as an Intermediate Format Generator
The direct generation of simulator-executable scenario files is inherently complex. To manage this complexity, the process is typically divided into multiple stages, with the outputs of each referred to in this study as intermediate formats. These formats acts as transitional representations that can be subsequently transformed into executable scenario files, as illustrated in step ③ of Figure 4. Unlike steps ① and ②, which primarily rely on either data or knowledge sources, step ③ integrates both.
A common application of LLMs in this context involves the generation of driving policies based on natural language inputs. In [60] and [61], user prompts are combined with environmental context and processed to derive driving strategies, which are subsequently used to inform either a driver model [60] or an auto-regressive trajectory generation model [61]. Similarly, Wei et al. [62] demonstrate the use of LLMs to interpret linguistic descriptions for trajectory planning, thereby guiding behavior of autonomous agents.
In addition, LLMs have been employed to synthesize various scenario elements. In [63], an approach is presented in which maps and vehicle assets are generated from textual descriptions. Waypoints are selected using a VLM, followed by trajectory generation using a diffusion model. In [64], critical corner cases are produced by integrating user language inputs, failure records, and scenario databases content. This methodology is further extended in [65], where scenario mutation, prior test feedback, and expert knowledge are incorporated to enhance scenario diversity. Similarly, in [66], descriptive inputs and datasets are processed to generate scenario configurations and associated parameter sets.
[…]

### v2

@@ II-A3 Scenario Abstraction Level
According to [menzel2018scenarios, asam2024openscenarioDSL], scenarios are organized into four levels of abstraction (see Figure 2): (1) Functional scenarios represent high-level traffic situations described in natural language using domain-specific terms. (2) Abstract scenarios are defined by transforming natural language descriptions into formats suitable for machine interpretation, often supported by ontologies [de2022towards]. (3) Logical scenarios define the parameter ranges of the state values used for scenario representation. (4) Concrete scenarios222A concrete scenario is not necessarily a test scenario. A test method requires a test scenario as input, which is derived from a concrete scenario but adapted to the ODD and testing requirements of the ADS. use specific state values to ensure their reproducibility and to enable test methods to execute the scenario.
II-A4 ODD
It comprises specific operating conditions within which an ADS is designed to function, enabling it to perform the dynamic driving task safely during a trip [ISO34502]. According to ISO/FDIS 34503 [ISO34503], ODD content is organized into three top-level attributes, as illustrated in Figure 3.
II-B Related Survey
Existing surveys provide overviews of LLMs, scenario-based testing, and ADSs, yet none has systematically examined their intersection. General surveys on LLMs cover architectures [Naveed.2024], training paradigms [hadi2023survey], and benchmark evolution [Zhang.2024], while others focus on reasoning efficiency [sui2025stop], prompt engineering [sahoo2024systematic], and privacy challenges [hu2023differentially, neel2023privacy, edemacu2024privacy]. In the ADS domain, scenario-based testing surveys provide taxonomies for verification methods [Nalic.2020, zhang2022finding], categorize scenario generation approaches as data-driven, adversarial, and knowledge-based [Ding.2023], review high-fidelity simulation platforms [Zhong.2021], and summarize verification frameworks and efficiency enhancements [Riedmaier.2020, Wishart.2020, Tang.2023]. Recent reviews [yang2023llm4drive, Zhou.2024b, Cui.2024, Fourati.2024, Li.2024d, Gan.2024] provide overviews of LLM-driven ADS research, highlighting perception, reasoning, and planning improvements enabled by multimodal extensions. Beyond ADSs, related surveys in software testing and urban systems [Braberman.2024, Zhou.2024, Xu.2024b] demonstrate the growing role of LLMs in intelligent verification and digital-twin applications. Moreover, large-scale European initiatives such as HEADSTART [sluis2021describing, op2021generation, wagner2020common] and SUNRISE [skoglund2025demonstrating, dokania2025implementing, vehicles7030100, de2024scenario, de2024coverage] have advanced the standardization and large-scale validation of scenario-based testing.
Beyond scenario-based testing, complementary lines of work in end-to-end closed-loop driving and benchmarking provide task definitions, evaluation criteria, and implementation baselines that can inform scenario generation and execution (e.g., [jia2023think, jia2024bench2drive, you2024bench2drive, jia2025drivetransformer, jia2023driveadapter, lu2024activead, yang2025drivemoe, yang2025raw2drive, li2024think2drive, wu2022trajectory]). In parallel, trajectory prediction and scene encoding methods support realism assessment and dynamic element modeling relevant to scenario fidelity (e.g., [jia2023hdgt, jia2024amp, jia2022multi, jia2023towards]). Additionally, human driving pattern mining and multimodal fusion contribute to richer behavior modeling and perception–planning interfaces (e.g., [jia2021ide, zhu2025flatfusion]).
Overall, prior works have remained domain-specific. The presented survey uniquely bridges these strands by structuring LLM-driven scenario-based testing research along testing phases and clarifying how LLMs contribute to each stage of ADS evaluation.
III LLM applications by phase
The scenario-based testing workflow, as illustrated in Figure 4, is initiated with scenario sources, which may include synthesized data, naturalistic driving measurements, or human expert knowledge. These sources are utilized to generate scenarios, which are subsequently organized and stored in a structured way within a scenario database. During the scenario selection phase, test scenarios are derived from the scenario database according to predefined criteria. The derived test scenarios are then executed on simulation platforms such as CARLA [pmlr-v78-dosovitskiy17a] to evaluate the performance of ADSs. Finally, test results are analyzed and reported in the ADS assessment phase.
In this section, the application of LLMs in scenario-based testing within different testing phases is examined, along with corresponding implementation details. An overview of the research structure for this section is presented in Figure 5. LLM applications are structured along five phases. In the scenario source phase (III-A), LLMs support data enrichment, labeling, and retrieval. In the scenario generation phase (III-B), they act as interpreters and format converters, enabling the transformation of NL or raw data into structured, standardized, and executable scenarios. In the scenario selection phase (III-C), LLMs contribute to realism assessment and anomaly detection. During test execution (III-D), they facilitate simulation setup automation and scenario optimization. Finally, in the ADS assessment phase (III-E), LLMs are applied to performance and intelligence-level evaluation.
III-A Scenario Source
Scenario-based testing of ADSs heavily relies on data, which serve as the basis for generating test scenarios. As a result, considerable attention has been directed toward scenario sources to support this process. As noted in [Nalic.2020, Riedmaier.2020], these sources can be classified into data-based or knowledge-based categories. Data may be obtained from real-world driving environments or synthetically generated using real-world contexts and expert input. Surveyed literature indicates that LLMs are primarily employed for tasks such as data enrichment, labeling, and retrieval. The implementation strategies associated with these applications are described below.
III-A1 Data Enrichment
The acquisition of real-world driving data often falls short of meeting the evolving demands of ADSs, primarily due to the high costs and time-intensive nature of data collection and annotation. As a result, data synthesis has become an essential component for ADS testing [song2023synthetic]. LLMs have been employed to generate driving trajectories based on NL instructions. In [Zhong.], LLMs enable bidirectional trajectory–language generation; the authors report a 50% reduction in the need for real data in overtaking prediction tasks; validation is limited to synthetic settings with simplified modality integration. In [Guo.2024], right-hand driving trajectories are synthesized by LLMs using left-hand driving data as a reference, the authors report improved cross-driving-side transfer on nuScenes [caesar2020nuscenes]; however, additional smoothing modules are required to ensure physically plausible trajectories. In [yang2025trajectoryllm], an LLM is used to translate brief descriptions of vehicle interactions into realistic driving behaviors and trajectories by incorporating human-like driving logic. Downstream predictors are improved when trained with synthetic data; however, performance is sensitive to multi-stage translation and to the annotation quality of driving logic.
Jia et al. [jia2023adriver] further expand data enrichment by combining an LLM with a video diffusion model to generate future driving scenes in video format. The authors report higher control accuracy on nuScenes [caesar2020nuscenes], but scalability is hindered by video quality degradation the lack of physical realism constraints in the generated data. Furthermore, LLMs have been adopted for ground truth generation; for example, Wei et al. [Wei.2024b] propose ChatSim, a framework that enables the creation of editable, photo-realistic 3D driving scenes through NL commands. Within this framework, LLMs are used to facilitate user interaction and to enable the integration of external digital assets. On WOD [sun2020scalability], the authors report higher realism and improved 3D detection, although background editing (e.g., weather) remains limited.
Overall, while LLM-driven data enrichment reduces the need for real data and broadens scenario diversity, most validations are limited to benchmark datasets and research simulators; cross-dataset/cross-platform replication and production-grade toolchain integration are rarely reported.
III-A2 Data Enrichment Through Hazard Analysis
HARA and STPA are originally developed for functional safety and system-level hazard analysis, respectively. These methods are also commonly applied within the scenario-based approach described in ISO 21448 (cf. [khatun2020scenario, khatun2025conceptual, zhang2022finding]) to systematically identify hazardous scenarios, gain insights into potential hazards (cf. [kramer2020identification, xing2021hazard, abdulkhaleq2018missing]) and guide the generation of high-risk scenario seeds (cf. [sun2025cascaded, zhang2024odd, khastgir2021systems]).
HARA automation using LLMs has been explored to improve the efficiency and accuracy of safety analysis in ADSs. At Qualcomm [Abbaspour.2024], LLMs demonstrate enhanced hazard identification capabilities, detecting 20% more hazards than conventional manual methods in an AEB case study. The relevance of hazards was ensured through an ISO 26262–aligned [ISO26262_2018] traceable chain verified by engineers, while the objective was comprehensive coverage rather than a fixed number, avoiding omission of high-ASIL risks. Volvo Cars researchers investigate LLMs to generate scenarios and malfunctions from functional descriptions using a general framework, without relying on standardized databases [Nouri.2024], thereby demonstrating their potential for domain-independent safety analysis. Validation is primarily via expert review; quantitative accuracy against manual baselines is not reported. A follow-up study [Nouri.2024b] applies LLMs to generate safety requirements and detect inconsistencies, which enhances efficiency in safety engineering while highlighting the need for expert validation to ensure accuracy.
STPA automation using LLMs has also been investigated. Charalampidou et al. [Charalampidou.2024] used ChatGPT-4 to generate loss scenarios for an unmanned aerial vehicle rescue system, the authors report a reduction in execution time (from weeks to hours), although the results require human validation. Similarly, Diemert et al. [diemert2023can] introduce a cooperative framework where LLMs assist in identifying hazards, though expert oversight is required due to inaccuracies reported, particularly for complex systems. Qi et al. [qi2025safety] extend this to automotive and energy systems, this study indicates that LLMs alone may be unreliable; certain human–LLM collaboration schemes outperform single-expert baselines on number/proportion of correct unsafe control actions, though results remain conservative and incomplete.
Most studies show that LLMs reduce manual effort in hazard analysis; however, outputs often exhibit inaccuracies and variability. Human inspection remains necessary as a guardrail. False‑positive/false‑negative rates, inter‑rater reliability, time‑to‑complete, and reproducibility are seldom reported. As found in [qi2025safety], human-LLM cooperation could be a promising solution. Moreover, failure cases are rarely presented or discussed, yet they could be valuable for improving LLMs’ hazard‑analysis capability.
III-A3 Data Labeling
Manual data annotation is both costly and labor-intensive, thereby necessitating automated solutions to enhance the scalability and efficiency of dataset development [song2023synthetic]. To address this challenge, LLMs have been utilized to partially substitute human effort in the annotation process. Chen et al. [chen2024automated] introduce the first benchmark for evaluating LVLMs in the context of self-driving corner cases. In their work, LLMs are employed to automatically generate annotations in JSON333https://www.json.org/json-en.html format, which were subsequently verified through manual inspection to ensure data accuracy and quality; quantitative metrics were not reported. While the authors mainly emphasize LLM-based annotation, the discussion of failure cases is limited. Nonetheless, expert oversight in such unlabeled or mislabeled cases remains crucial, as these situations may offer the most valuable insights for advancing scenario understanding.
III-A4 Data Retrieval
Effective scenario generation requires the retrieval of relevant information from large-scale, heterogeneous datasets, which often include multimodal data in varying formats and structures. This inherent complexity presents challenges for unified processing and has traditionally been managed through rule-based methods, such as SQL queries (cf. [babisch2023leveraging]), which demands significant manual effort. Recent studies have instead employed LLMs to enable more intuitive and flexible data retrieval, thereby reducing the reliance on predefined rules and manual querying.
Video retrieval efficiency has been demonstrated to improve through the application of MLLMs. In the study by Knapp et al. [knapp2024data], MLLMs are integrated with a vector database to enable NL querying of driving logs. Scenario descriptions were generated from sensor data and video content, allowing users to navigate large-scale datasets without relying on SQL-based queries. The authors report quantitative experiments and an engineer survey, demonstrating the method’s practical utility. However, the approach remains constrained by models’ limited context size and single-frame inputs, preventing full signal utilization and reducing awareness of temporal dynamics.
Image retrieval efficiency has also been improved through the application of LVLMs. In [Sohn.2024], three LVLMs are employed to perform image retrieval based on NL queries. Their performance is evaluated using precision-based metrics on BDD100K [yu2020bdd100k]. The models achieve high precision and capture details such as color and position, but show weakness in temporal information like maneuvers and may hallucinate in motion-related scenarios.
Additionally, Rigoll et al. [Rigoll.2024] propose an annotation-free, object-level image retrieval approach for AD datasets by combining panoptic segmentation with a VLM to support NL queries. While this approach enhances data accessibility without manual labeling, it is limited to single-object recognition and performs poorly in complex, dynamic scenes involving multiple objects. To overcome these limitations, Tang et al. [Tang.2024] incorporate LLMs and BEV representations to enhance text-to-scene retrieval. On the nuScenes-Retrieval dataset, an extension of nuScenes [caesar2020nuscenes] with enriched textual annotations, the authors report state-of-the-art performance, outperforming baselines such as CLIP-ViT-Base [radford2021learning], SigLIP-Base [zhai2023sigmoid], and EVA02-Base [fang2024eva]. However, challenges remain in handling multiple agents and long, complex textual descriptions.
Most above studies are evaluated on narrow benchmarks and research-oriented settings. Despite support from human studies and quantitative tests, they are typically constrained by context-window limits and frame-level inputs, which hinder effective use of long-horizon multimodal signals. Image retrieval often degrades in complex multi-agent or occluded scenes, and some methods are restricted to single-object queries. Even when reporting state-of-the-art results on nuScenes‑Retrieval, works seldom disclose cross-dataset robustness, retrieval latency/throughput, index update cost and memory footprint, or key reproducibility details (model version, prompts, sampling parameters, random seeds), limiting assessments of practical deployability and transferability.
III-B Scenario Generation
This section reviews previous studies that have employed LLMs for scenario generation. To enable execution within simulation platforms such as CARLA [pmlr-v78-dosovitskiy17a], the generated scenarios must be converted into standardized formats, such as XOSC [ASAM_OpenSCENARIO_2.0.0]. Given the inherent complexity of scenario generation, the process is typically decomposed into several sub-steps. When LLMs are utilized, their contributions are generally restricted to specific stages of the workflow rather than spanning the entire process. An overview of the functional roles played by LLMs throughout the scenario generation pipelines is illustrated in Figure 6, and the subsequent discussion is organized accordingly.
III-B1 LLM as a Human-Machine Interface
In this subsection, LLMs are employed to translate NL inputs provided by the user into structured information that facilitates downstream scenario generation, as illustrated in step ① of Figure 6. Specifically, LLMs have been used to interpret user input into structured scenario representations [10588843, Li.2024, Tan.2023, cai2025text2scenario, jiang2024scenediffuser], loss function formulations [Liu.2024b, ZiyuanZhong.2023], and executable code [Xia., Zhao.2024b, Nguyen.2024]. The primary strategies adopted in these studies are schematically summarized in Figure 7.
Structured scenario representations derived from user inputs have been employed in various studies to support scenario generation. In [Liu.2024b], such structured information is employed to create executable files through a diffusion-based model. Similarly, in [10588843] and [Li.2024], it is used to trigger Python scripts for automated file generation. Building upon [10588843], Zhou et al. [Zhou.21022025] evaluate the performance of six different LLMs in interpreting motorway functional scenarios, thereby offering a benchmark for model selection. In [Tan.2023], structured representations condition a transformer model trained on the WOD[sun2020scalability] for learning realistic traffic distributions to generate vehicle trajectories. This approach outperforms baselines TrafficGen[feng2023trafficgen] and MotionCLIP[tevet2022motionclip] with improved realism and controllability, yet remains constrained by the lack of lane-level topology. In [cai2025text2scenario], LLMs convert NL inputs into scenario elements, which are subsequently transformed into simulation-executable files. It demonstrates notable efficiency gains and competitive accuracy across multiple AD stacks. However, it remains constrained by output instability, reliance on predefined knowledge bases, and limited format generalization. A unified diffusion-based framework is introduced in [jiang2024scenediffuser], where LLMs generate Proto444https://protobuf.dev/ constraints from user input to enable language-driven control of scene initialization and closed-loop simulation.
Loss functions have also been generated from NL inputs using LLMs. In[Liu.2024b], LLMs generate cost functions via a CoT module to guide a diffusion model, iterative refinement to improve trajectory quality. On WOD [sun2020scalability], the method enables controllable scene simulation with high success rates and zero-shot accident generation, while fundamental LLM misinterpretations remain uncorrected. In [ZiyuanZhong.2023], a loss function is derived from user-provided descriptions and integrated into a diffusion model to generate multi-agent traffic trajectories. On nuScenes dataset[caesar2020nuscenes], it surpass CTG[zhong2022guided] and BITS[xu2022bits] in realism and controllability but remains limited by map-intensive instructions, error-prone loss functions, and slow generation.
Code generation from NL inputs is further demonstrated in [Xia., Zhao.2024b]. In [Xia.], the generated code, which includes interaction, vehicle, and map modules, is utilized by a transformer to synthesize vehicle trajectories. This approach achieves lower displacement errors and more realistic interaction behaviors on WOMD [ettinger2021large], compared against TrafficGen [feng2023trafficgen] and LCTGen [Tan.2023]; but it is limited to generating vehicle trajectories, without support for other traffic participants or map dynamics. In [Zhao.2024b], code-generated trajectories conditions a map generator and a video model to synthesize diverse driving videos; on nuScenes [caesar2020nuscenes], it improves visual fidelity, temporal coherence, and downstream perception metrics. Nguyen et al. [Nguyen.2024] likewise report improved behavioral diversity on the same dataset, though unintended behaviors remain.
III-B2 LLM as a Data Interpreter
This subsection reviews studies in which LLMs have been employed to interpret various data sources into NL representations. This process is illustrated in step ② of Figure 6. Scenario-relevant information has been extracted from diverse inputs, including accident reports [Guo.2024b], domain-specific knowledge [Tang.2023c], and naturalistic driving data [mei2025llm, xu2025chatbev, tian2024enhancing].
Accident reports have been interpreted by LLMs to support scenario reconstruction. In Guo et al. [Guo.2024b], NHTSA accident reports [NHTSA2024NMVCCS] are parsed into a structured JSON format, which is subsequently processed by a constraint solver to generate waypoints based on road geometries. The approach reports high information extraction accuracy and generalizes well at intersections, while testing on an industrial ADS exposed multiple safety violations. Additionally, domain-specific knowledge has been extracted using LLMs to aid scenario generation. Tang et al. [Tang.2023c] transform technical documents into structured ontological elements for automated ontology construction. However, output quality is variable and prone to semantic errors, requiring human oversight for reliability
Naturalistic driving data have also been utilized to support scenario generation through LLM-based interpretation. Mei et al. [mei2025llm] employ LLMs to identify the most threatening attacker in each scene from the WOMD[ettinger2021large], enabling the creation of high-risk scenarios for improving the testing and training of ADSs; in MetaDrive [9829243], the approach reports higher attack success rates and improved ADS safety, though efficiency is limited by computational cost. In [xu2025chatbev], LLMs interpret BEV maps derived from the nuPlan dataset [caesar2021nuplan] to guide a diffusion model for trajectory generation. It outperforms deep learning and zero-shot VLM on scene understanding and improved trajectory fidelity, but remains restricted to static, vehicle-centric scenarios without dynamic scenes or pedestrian activities. Tian et al. [tian2024enhancing] combine RL with an LLM in HighwayEnv [Leurent_HighwayEnv]: by analyzing failure modes and risk metrics, the LLM suggests environment modifications that increase scenario diversity and policy challenge; compared with PPO, training loss decreases, rewards increase, and collisions are reduced, while validation is confined to HighwayEnv and requires broader testing. For online generation, Mei et al. [mei2025seeking] use retrieval-augmented LLMs to infer dangerous behaviors and synthesize adversarial trajectories based on historical states. It demonstrates lower minimum TTC, and higher collision rate in WOMD [ettinger2021large] compared to transformer-based approaches.
III-B3 LLM as an Intermediate Format Generator
The direct generation of simulator-executable scenario files is inherently complex. To manage this complexity, the process is typically divided into multiple stages, with the outputs of each referred to in this study as intermediate formats. These formats act as transitional representations that can be subsequently transformed into executable scenario files, as illustrated in step ③ of Figure 6. Unlike steps ① and ②, which primarily rely on either data or knowledge sources, step ③ integrates both.
A common application of LLMs in this context involves the generation of driving policies based on NL inputs. In [zhou2024humansim] and [Tan.2024], user prompts are combined with environmental context and processed to derive driving strategies, which are subsequently used to inform either a driver model [zhou2024humansim] or an auto-regressive trajectory generation model [Tan.2024]. Similarly, Wei et al. [Wei.2024] demonstrate the use of LLMs to interpret linguistic descriptions for trajectory planning, thereby guiding behavior of autonomous agents.
In addition, LLMs have been employed to synthesize various scenario elements. DriveGen [zhang2025drivegeninfinitediversetraffic] generates maps and vehicle assets from textual descriptions; a VLM selects waypoints and a diffusion model produces trajectories. It yields diverse, realistic scenarios that benefit downstream AV algorithms, though VLM-based waypoint selection introduces small trajectory inconsistency compared with BITS[xu2022bits]. In [Chang.2024], critical corner cases are produced by integrating user language inputs, failure records, and scenario databases content. This approach expands the diversity of rare and risky scenarios, though its evaluation is limited to highway data. In [Xu.2024], scenario mutation, prior test feedback, and expert knowledge are incorporated into a “generate-test-feedback” loop that reveals policy failures but remains constrained by LLM’s difficulty with with precise numerical adjustments and suffers from API-call latency. Similarly, Li et al. [Li.2024c] convert descriptive inputs and datasets into scenario configurations and parameter sets, achieving high accuracy in mixed vehicle-pedestrians tasks and more human-like motion dynamics, yet the current scope is limited to vehicles and pedestrians.
[…]